OpenStudy (anonymous):

What are the missing y-values in the table below? y = x - 2 x y 3 ? 6 ? 8 ? 10 ?

OpenStudy (anonymous):

*Stares blankly*

OpenStudy (anonymous):

how do you do this?

OpenStudy (anonymous):

you just need to plug each x value into the equation y = x-2 Here, I'll do the first y = x-2 x = 3 y= 3-2 = 1

OpenStudy (anonymous):

do the same thing for the other x values

OpenStudy (anonymous):

oh ok thanks

OpenStudy (anonymous):

sure thing. good luck!